4 Two sheet metal parts, A and B, are produced separately, each requiring two press-working operations. Part A is routed through operations 1 and 2, and part B is routed through oper- ations 3 and 4.The two parts are then joined in a welding step (operation 5), and the assem- bly is routed to an electroplating operation (operation 6). The six operations have the following fraction defect rates: qi-0.05, q,-0.15.q,-0.10, q,-0.20, q,-0.13, q.-0.08. If the desired final quantity of assemblies is 100,000 units, how many starting units of parts A and B will be required? There is no inspection or separation of defective units until after the final process, so defective units and good units are processed together through all pro- duction processes.
Answer #1

Required: output of Q6 100,000 pc o, (A) 2 q0.05 q-0.15 100,000 0.13 q 0.08 o, (B) 3 4 q 0.10 q0.20= Q5(1 – 0.08), Q5= 100,000/(0.92) = 108,696 pc

With q5= 0.13, input to Q5= 108,696/(1 – 0.13) = 124,938 pc

Starting quantity Qo(A) = Qo(B) = 124,938/{(1-0.05)(1-0.15)(1-0.10)(1-0.20)}

= 124,938/0.5814 = 214,891 pc of A and 214,891 pc of B

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW

PLEASE RATE MY ANSWER BY HITTING ??

THANK YOU